need help Bobbys grandfather is 6 years older than 7 times bobbys age.the difference in their ages is 72years.how old is bobby and his grandfather
Answer by stanbon(75887)

Bobbys grandfather is 6 years older than 7 times bobbys age.the difference in their ages is 72years.how old is bobby and his grandfather
--------------------
Equations:
G = 7B+6
G - B = 72
------------------
Substitute for "G" and solve for "B":
7B+6 - B = 72
---
6B = 66
---
B = 11 yrs old (Bobby's age now)
----
Solve for "G":
G = 7B+6
G = 7*11+6 = 77+6 = 83 yrs old (Grandpa's age now)
======================================================
Cheers,
Stan H.
===============